# Stark localization near Aubry-André criticality

### v1

@@ Abstract
In this work we investigate the Stark localization near the Aubry-André (AA) critical point. We study system-dependent parameters, such as localization length, inverse participation ratio (IPR), and energy gap between the ground and first excited state, for characterizing the localization-delocalization transition. We show that the scaling exponents possessed by these key descriptors of localization are quite different from that of a pure AA model or Stark model. Near the critical point of the AA model, inducing Stark field of strength $h$, the localization length $\zeta$ scales as $\zeta\propto h^{-\nu}$ with $\nu\approx 0.29$ which is different than both the pure AA model ($\nu=1$) and Stark model ($\nu\approx 0.33$). The IPR scales as IPR $\propto h^{s}$ with $s\approx 0.096$ which is again significantly different than both the pure AA model ($s\approx 0.33$) and Stark model ($s\approx 0.33$). The energy gap, $\Delta$, scales as $E\propto h^{\nu z}$, where $z\approx 2.37$ which is however same as the pure AA model.
@@ I Introduction
Randomness is an inherent property of nature whose influence on any physical system often defies our intuition and yields non-trivial effects. Consequently, disordered systems have long been a focal point of significant research endeavors Binder86 ; Belitz05 ; Das08 ; Alloul09 ; Konar22 ; Ghosh21 . Within the realm of quantum mechanics, the interplay of disorder with fundamental principles gives rise to a myriad of noteworthy phenomena, including the emergence of novel quantum phases, order-from-disorder phenomena Aizenman89 ; Wehr06 ; Bera14 ; Bera16 ; Bera17 , the enigmatic high-$T_{c}$ superconductivity Auerbach94 , and the phenomenon of localization Anderson58 ; Abrahams79 ; Evers08 ; Nandkishore15 ; Abanin19 ; Aitijhaya23 . Among these phenomena, disorder-induced Anderson localization occupies a prominent position Yamilov23 ; Lagendijk09 . This well-established concept elucidates how the presence of randomness can hinder the diffusion of waves within a system, showcasing its pivotal role in modern condensed matter physics. In Anderson model the disorder is truly random and uncorrelated. Whereas some disorders are not random, rather there is some correlation. Quasi-periodicity is the candidate that qualifies for this correlated disorder.
@@
The AA model is a tight binding model with nearest neighbour tunnelling in the presence of sinusoidal quasi-periodic potential. When the strength of the potential becomes twice the tunneling parameter, in the thermodynamic limit, the system experiences a phase transition from a delocalized phase to a localized phase. On the other hand, in a tight binding model, if a gradient field is induced across the lattice making the on-site energies off-resonant, the tunnelling rate gets suppressed resulting in the localization of wavefunction in space in absence of disorder. In the thermodynamic limit, this localization, known as Stark localization Schulz19 , takes place at zero field limit. A substantial amount of research has been conducted for understanding Stark localization phenomena in context of single-particle Wannier60 ; Fukuyama73 ; Holthaus95 ; Kolovsky03 ; Kolovsky08 ; Kolovsky13 ; Nieuwenburg19 and many-body Jiang23 ; Zhang21 ; Yao21 ; Wei22 ; Doggen21 systems, and for finding applications Morong21 ; Preiss15 ; Karamlou22 ; He23 ; Kohlert21 by exploiting it.
@@
All these are quantum phase transitions that manifest at zero temperature when the control parameter of a quantum system’s Hamiltonian is finely tuned to a critical value, referred to as the quantum critical point (QCP). At this juncture, the system exhibits scaling and universality Hikami81 ; Altland97 ; Wang21 ; Luo21 ; LuoR22 ; Luo22 ; Bu22 , meaning that the equilibrium properties of certain physical observables near the QCP can be succinctly described by a handful of critical exponents. Extracting these critical exponents necessitates an understanding of various physical quantities  Mishra16 ; Sadhukhan15 ; Sadhukhan16 , such as the localization length, Inverse Participation Ratio (IPR), energy gap, and fidelity susceptibility etc. These quantities provide essential insights into the behavior of the system near the QCP and facilitate the characterization of its critical behavior. In this work, we consider a single particle tight-binding model with nearest-neighbour hopping. The system is prepared at the QCP of the AA model. Then a Stark potential is introduced for studying the scaling properties of localization length, IPR, and energy gap between the ground and the first excited state.
@@
Following an introductory overview in Section I, Section II presents the model of our system. In Section III, defining the characteristic quantities relevant to the localization-delocalization transition, the behavior of the corresponding scaling exponents at the criticality of the AA model is studied via the cost function approach. In Section IV the same scaling exponents are computed in the vicinity of the critical point of the AA model. In Section V, we explore the evolution of the scaling exponent of IPR as we move away from the critical region of the AA model and additionally, we discuss how the corresponding structures of the wave function change. Then in Section VI, we introduce a hybrid scaling exponent that is related to the other scaling exponents of the pure AA and Stark model. Finally, a conclusion is given in Section VII.
@@ II Model
$\hat{H}=$ $-\;J$ $\sum_{i}^{L}(\hat{c}^{\dagger}_{i}\hat{c}_{i+1}+\mathrm{h.c.})+h\sum_{i}^{L}i\hat{c}^{\dagger}_{i}\hat{c}_{i}$ (1)
$+$ $(2J+\delta)\sum_{i}^{L}\cos[2\pi(i\omega+\phi)]\hat{c}^{\dagger}_{i}\hat{c}_{i}$
where, $\hat{c}^{\dagger}_{i}(\hat{c}_{i})$ is the fermionic creation (annihilation) operator at the $i^{\mathrm{th}}$ site, $(2J+\delta)$ determines the amplitude of the quasiperiodic modulation, $\omega$ is an irrational number has a special form of $\omega=F_{n}/F_{n+1}$ where $F_{n}$ are the Fibonacci numbers and system size $L$ has been chosen as $L=F_{n+1}$, $\phi$ is a random number resulting in some random phase of quasiperiodic potential, drawn from a uniform distribution in $[0,1)$, and $h$ quantifies the magnitude of the Stark field. The boundary condition used in this work is the open boundary condition (OBC). In the thermodynamic limit (i.e., $L\to\infty$) for the absence of the second term, i.e., $h=0$, one finds that all the eigenstates are localized for $\delta>0$ and all the eigenstates are delocalized for $\delta<0$. On the other hand, without the last term i.e., $\delta=-2J$, in the thermodynamic limit, the system is localized for $h\to 0$.
@@ III Scaling Analysis at the AA Criticality
We start our study of the behaviour of the scaling exponents right at the AA criticality, i.e., $\delta=0$. In a finite system, at this point, the wavefunction of the particle is neither localized nor delocalized, rather there happens to be a continuous phase transition. To quantify the localized or delocalized nature of the wavefunction, we use the observable localization length, $\zeta$, defined as,
@@
where $i$ denotes the lattice site, $p_{i}$ is the single-particle probability density at site $i$ and $i_{c}$ is the localization center having the expression $i_{c}=\sum_{i=1}^{L}ip_{i}$. Expanding the $n^{\mathrm{th}}$ normalised eigenstate of the system, $|\psi^{n}\rangle$, in terms of the single-particle computational basis, $|i\rangle$, such that $|\psi_{n}\rangle=\sum_{i}c^{(i)}_{n}|i\rangle$, $p_{i}$ is given by $p_{i}=|\langle i|\psi^{n}\rangle|^{2}=|c^{(i)}_{n}|^{2}$.
Figure 1: Scaling Analysis at $\delta=0$: (a1) presents localization length $\zeta$ of the ground state against Stark field strength $h$ for different system sizes, $L=144$ (blue circle), $L=233$ (violet square), $L=377$ (indigo diamond), $L=610$ (maroon triangle up) and $L=987$ (orange triangle left). (a2) shows a collapse plot for the localization length with the value of scaling exponent $\nu=0.29$. (b1) displays the IPR versus $h$ for various values of $L$ which is aforementioned in (a1). (b2) demonstrates the collapse plot of IPR using the value of scaling exponent $s=0.096$ and $\nu=0.29$. (c1) shows the energy gap $\Delta E$ between the ground state and the first excited state as a function of $h$ for different $L$ values as mentioned previously. (c2) demonstrates the collapse plot of $\Delta E$ using the scaling exponent $z=2.37$. we have taken $\phi$ average for 5000 random choices in the range $[0,1)$.
If $g$ is the control parameter in a certain localization-delocalization transition and $g_{c}$ is the critical point then in the thermodynamic limit, near this criticality, the localization length scales as,
@@
where $f_{1}[\cdot]$ is an arbitrary function. This scaling ansatz is confirmed through the methodology of data collapse. The concept behind data collapse is as follows: We collect the data $\zeta/L$ versus $hL^{1/\nu}$ for various system sizes $L$, with a specific value of $\nu$, for which all the datasets collapse perfectly. Initially, by some trial and error, we manually find out a rough range of the scaling exponent for which the data collapse is quite well. Furthermore, for a better estimation of the scaling exponent we employ the cost function approach Suntajs20 , where the cost function $C_{Q}$ is defined as,
In Fig. 1(a1), we depicted $\zeta$ against $h$ for various values of $L$. This figure illustrates that for finite-sized systems, an initial flat region exists wherein the corresponding wave function exhibits an extended nature, indicating that the ground state of the Hamiltonian for $\delta=0$ is equally probable across this phase. Beyond a certain threshold of $h$, the localization length exhibits weak dependence on the system size, leading to the localization of the wavefunction. To derive the scaling exponent $\nu$ from Eq. (3) with $|g-g_{c}|=h$, we demonstrate a collapse plot in Fig. 1(a2) utilizing the function specified previously in Eq. (5). The best collapse is obtained for $\nu=0.29(1)$. The cost function approach further confirms this optimal value of the scaling exponent $\nu$.
@@
The variation in the value of IPR with the increase in $h$ for different system sizes $L$ is depicted in Fig. 1(b1). In this figure, upto a certain region of $h$, the IPR values appear flat, indicating the presence of a delocalized phase. Beyond this region, the IPR becomes independent of the system size, characterizing the localized phase. We employ the cost function approach which suggests $s=0.096(1)$. Fig. 1(b2) presents the best collapse plot of the IPR using $s=0.096$. Noticeably, in this case, $s/\nu\approx 0.33$ which is the same as the pure AA model demonstrating the fact that at the criticality the scaling of IPR with the system size is the same for both of these two models.
@@
Fig. 1(c1) shows, the energy gap $\Delta E$ against the Stark field strength $h$ for different system sizes. In this case, also there is an initial flat region showing a very narrow energy gap in the delocalized phase. Beyond a certain strength of Stark field $h$, $\Delta E$ becomes independent of the system size and the energy gap gets increased with an increase in the field strength. The scaling exponent $z$ is determined through data collapse yielding a critical exponent of $z=2.37(1)$. It is also the same as the pure AA model. The collapse plot corresponding to $\Delta E$ is presented in Fig. 1(c2).
Figure 4: (a1) presents $s/\nu$ for different values of $\delta$. For $\delta=0$ the $s/\nu$ value is $0.33$ which is the same as that of the pure AA model and this value continuously increases to $1$, which is the same as the pure Stark model, as one moves away from the AA criticality with negative $\delta$ values. (a2) presents the plot of $z$ for different values of $\delta$. In this case also, there is a continuous change in the value of $s/\nu$ from 2.37, which is the same as the pure AA model, to 2, which is for the pure Stark model, as one moves away from the criticality.

### v2

@@ Abstract
In this work, we investigate the Stark localization near the Aubry-André (AA) critical point. We perform careful studies for reporting system-dependent parameters, such as localization length, inverse participation ratio (IPR), and energy gap between the ground and first excited state, for characterizing the localization-delocalization transition. We show that the scaling exponents possessed by these key descriptors of localization are quite different from that of a pure AA model or Stark model. Near the critical point of the AA model, in the presence of Stark field of strength $h$, the localization length $\zeta$ scales as $\zeta\propto h^{-\nu}$ with $\nu\approx 0.29$ which is different than both the pure AA model ($\nu=1$) and Stark model ($\nu\approx 0.33$). The IPR in this case scales as IPR $\propto h^{s}$ with $s\approx 0.096$ which is again significantly different than both the pure AA model ($s\approx 0.33$) and Stark model ($s\approx 0.33$). The energy gap, $\Delta$, scales as $E\propto h^{\nu z}$, where $z\approx 2.37$ which is however same as the pure AA model. Finally, we discuss how invoking a criticality inducing additional control parameter may help in designing better many-body quantum sensors. Quantum critical sensors exploit the venerability of the wavefunction near the quantum critical point against small parameter shifts. By incorporating a control parameter in the form of the quasi-periodic field, i.e., the AA potential, we show a significant advantage can be drawn in estimating an unknown parameter, which is considered here to be the Stark weak field strength, with high precision. We compute and extract the scaling exponent associated with the Quantum Fisher Information (QFI), $F_{Q}$, which provides a bound on the uncertainty associated with measuring an unknown parameter. We demonstrate that by a control field in the form of the quasiperiodic potential, and setting it at AA criticality, $F_{Q}$ scales with the system size as, $F_{Q}\propto L^{\beta}$, where $\beta\approx 6.7$, which is greater in comparison to the pure Stark case, for which $\beta$ is known to be $5.8$. We, thus, pitch a generic idea of invoking quantum criticality inducing additional control parameters for obtaining enhanced quantum advantage in many-body critical sensors.
@@ I Introduction
Randomness is an inherent property of nature whose influence on any physical system often defies our intuition and yields non-trivial effects. Consequently, disordered systems have long been a focal point of significant research endeavors Binder86 ; Belitz05 ; Das08 ; Alloul09 ; Konar22 ; Ghosh21 ; Aitijhaya23 ; roy20 . Within the realm of quantum mechanics, the interplay of disorder with fundamental principles gives rise to a myriad of noteworthy phenomena, including the emergence of novel quantum phases, order-from-disorder phenomena Aizenman89 ; Wehr06 ; Bera14 ; Bera16 ; Bera17 ; Mishra16 ; Sadhukhan15 ; Sadhukhan16 , the enigmatic high-$T_{c}$ superconductivity Auerbach94 , and the phenomenon of localization Anderson58 ; Abrahams79 ; Evers08 ; Nandkishore15 ; Abanin19 . Among these phenomena, disorder-induced Anderson localization occupies a prominent position Yamilov23 ; Lagendijk09 . This well-established concept elucidates how the presence of randomness can hinder the diffusion of waves within a system, showcasing its pivotal role in modern condensed matter physics. In Anderson’s model the disorder is truly random and uncorrelated. Whereas some disorders are not random, rather there is some correlation. Quasi-periodicity is the candidate that qualifies for this correlated disorder.
@@
The AA model is a tight binding model with nearest neighbor tunneling in the presence of sinusoidal quasi-periodic potential. When the strength of the potential becomes twice the tunneling parameter, in the thermodynamic limit, the system experiences a phase transition from a delocalized phase to a localized phase. On the other hand, in a tight binding model, if a gradient field is induced across the lattice making the on-site energies off-resonant, the tunneling rate gets suppressed resulting in the localization of wavefunction in space in the absence of disorder. In the thermodynamic limit, this localization, known as Stark localization Schulz19 , takes place at zero field limit. A substantial amount of research has been conducted for understanding Stark localization phenomena in context of single-particle Wannier60 ; Fukuyama73 ; Holthaus95 ; Kolovsky03 ; Kolovsky08 ; Kolovsky13 ; Nieuwenburg19 and many-body Jiang23 ; Zhang21 ; Yao21 ; Wei22 ; Doggen21 systems, and for finding applications Morong21 ; Preiss15 ; Karamlou22 ; He23 ; Kohlert21 by exploiting it.
Figure 1: A schematic figure depicting the nature of localization in the presence of the dual potentials. In the thermodynamic limit, the system is undet the influence of the joint criticalities at $h\to 0$ and $\delta=0$ due to the Stark potential and the AA potential, respectively (black circle). The system is delocalized for $|\delta|>0$ and $h=0$ (red line). It enters into the localized phase in the limit $h\to 0$, irrespective of the value $\delta$. However, the finite value of $\delta$ may significantly influence the nature of localization when compared to the pure Stark case, and it is then termed as AAS critical localization (yellow region). For large values of $h$ and away from $\delta=0$, the AA potential has negligible influence on localization, and it becomes Stark dominated, as one may expect. In the $h\to 0$ limit, this can by shown by comparing the scaling exponent associated with the localization transition with that of the pure Stark model. For non-zero $h$, we monitor the fidelity of the ground state wavefunction with the system under joint potentials with the pure Stark case. The fidelity is expected to be high in the Stark dominated localization (blue region).
@@
All these are quantum phase transitions that manifest at zero temperature when the control parameter of a quantum system’s Hamiltonian is finely tuned to a critical value, referred to as the quantum critical point (QCP). At this juncture, the system exhibits scaling and universality Hikami81 ; Altland97 ; Wang21 ; Luo21 ; LuoR22 ; Luo22 ; Bu22 , meaning that the equilibrium properties of certain physical observables near the QCP can be succinctly described by a handful of critical exponents. Extracting these critical exponents necessitates an understanding of various physical quantities, such as the localization length, Inverse Participation Ratio (IPR), energy gap, fidelity susceptibility, etc. These quantities provide essential insights into the behavior of the system near the QCP and facilitate the characterization of its critical behavior. In this work, we consider a single particle tight-binding model with nearest-neighbor hopping. Particularly, quantum criticality in the disordered AA model has been studied recently and a hybrid scaling theory has been developed Bu22 . In a similar spirit, we ask the following question: What is the nature of the quantum criticality of the AA system in the presence of an additional Stark potential? The Stark potential is always localizing. However, the interplay of these potentials ensures that the system has a rich quantum of critical behavior. The system is first prepared at the QCP of the AA model. Then a Stark potential is introduced for studying the scaling properties of localization length, IPR, and energy gap between the ground and the first excited state. We perform several case studies that include scaling studies when the AAH field is set (i) right at the AAH criticality, (ii) near the AAH criticality, (iii) far from the AAH criticality, and (iv) a case, where the system size, $L$, is kept fixed and different AAH field strengths are considered in order a perform a hybrid scaling analysis. The modified nature of the criticality of the system under the joint presence of the localizing potentials is completely captured in the associated scaling exponents. We perform a careful error analysis for reporting the critical exponents obtained from the scaling analysis. The scaling exponents are estimated to have uncertainties of the order of 1%. It’s worth mentioning that the scaling theory in the AAS model can be formulated for the range of the AA potential that corresponds to the delocalized phase up to the AA criticality in a pure AA model. Beyond the AA criticality where the AA potential induces a finite localization in the pure AA model, any kind of scaling analysis in the AAS model fails. One may understand that for the regime of the AA potential, where scaling theory can be properly devised, the scaling exponents in the AAS system is strongly modified when the AA potential is at or near the AA criticality. It will be shown that far away from the AA criticality, the AA potential exerts negligible influence and the physics is dominated by the Stark potential. As a result, the scaling exponent describing the criticality of the AAS system becomes identical to the pure Stark system. Then the ground state wave function of the AAS system has high fidelity with the ground state of the pure Stark model. This fidelity based analysis is extended for identifying the parametric regime where localization properties are basically dominated by the Stark potential alone.
Apart from the fundamental interest, we show that the system under study has important applications in quantum metrology QM006 . Recently, there are a lot of efforts towards designing critical quantum sensors - a new kind of many-body quantum sensing device. Several of these works have recognized quantum phase transitions, such as first- and second-order quantum phase transition, and localization transition, as useful quantum resources that can be exploited for engineering the critical quantum sensors He23 ; Rams18 ; Sahoo24 ; Sahoo24stark ; Mondal24 . Estimating the strength of an unknown weak field with high precision is a formidable task. A recent work has proposed a weak-field sensing device by employing Stark weak field inducing localization He23 . Moreover, the authors of this work also demonstrated in a parallel work that localization-delocation transition in the AA can be an excellent resource for designing many-body quantum sensors that can estimate an unknown field with high accuracy Sahoo24 . Performance of these quantum sensing devices are marked via computation of Quantum Fisher Information (QFI), $F_{Q}$, that provides a bound in uncertainty associated with the measurement of the unknown parameter. It is known that the QFI of a quantum critical sensor is controlled by the scaling exponent associated with the correlation length (in case of second-order quantum phase transition) or localization length (in case of localization transition) He23 ; Sahoo24 . Now, given the fact that the interplay of these Stark and AA potentials modifies quantum critical behavior and the scaling properties of localization length, a natural question arises if it is possible to design a better Stark weak field sensor by invoking an additional control parameter in the form of the AA potential, and, in particular, by setting it near the AA criticality. The answer is affirmative. We present an analysis in this work in support of this claim.
@@
Following an introductory overview in Section I, Section II presents the model of our system. In Section III, defining the characteristic quantities relevant to the localization-delocalization transition, the behavior of the corresponding scaling exponents at the criticality of the AA model is studied via the cost function approach. In Section IV the same scaling exponents are computed in the vicinity of the critical point of the AA model. In Section V, we explore the evolution of the scaling exponent of IPR as we move away from the critical region of the AA model and additionally, we discuss how the corresponding structures of the wave function change. Then in Section VI, we introduce a hybrid scaling exponent that is related to the other scaling exponents of the pure AA and Stark model. Finally, we discuss how the model under study can be utilized for designing better quantum sensors for precise estimation of the Stark weak field amplitude in Section VII. A conclusion is given in Section VIII.
@@ II Model
$\hat{H}=$ $-\;J$ $\sum_{i}^{L-1}(\hat{c}^{\dagger}_{i}\hat{c}_{i+1}+\mathrm{h.c.})+h\sum_{i}^{L-1}i\hat{c}^{\dagger}_{i}\hat{c}_{i}$ (1)
$+$ $(2J+\delta)\sum_{i}^{L-1}\cos[2\pi(i\omega+\phi)]\hat{c}^{\dagger}_{i}\hat{c}_{i}$
where, $\hat{c}^{\dagger}_{i}(\hat{c}_{i})$ is the fermionic creation (annihilation) operator at the $i^{\mathrm{th}}$ site, $(2J+\delta)$ determines the amplitude of the quasiperiodic modulation that is controlled by an irrational frequency $\omega$. It is well known in the literature that proper scaling with system size $L$ appears if it is chosen from the Fibonacci series, $L=F_{n+1}$. For finite size systems, $\omega$ is approximated as $\omega=F_{n}/F_{n+1}$. Here $F_{n}$ and $F_{n+1}$ are two consecutive Fibonacci numbers. In the thermodynamic limit, $n\to\infty$, it has the property $\omega=\text{lim}_{n_{\to\infty}}F_{n}/F_{n+1}=(\sqrt{5}-1)/2$, which is known as the golden ratio. $\phi$ is a random number resulting in some random phase of quasiperiodic potential, drawn from a uniform distribution in $[0,1]$, and $h$ quantifies the magnitude of the Stark field. The boundary condition used in this work is the open boundary condition (OBC). In the thermodynamic limit (i.e., $L\to\infty$) for the absence of the second term, i.e., $h=0$, one finds that all the eigenstates are localized for $\delta>0$ and all the eigenstates are delocalized for $\delta<0$. On the other hand, without the last term i.e., $\delta=-2J$, in the thermodynamic limit, the system is localized for $h\to 0$. For the rest of the work, we fix $J$ at unity, i.e., $J=1$, for convenience.
Before we delve deeper into the details, we present a schematic diagram as a precursor of the results in Fig. 1 of the Aubery-André-Stark (AAS) model in the parameter regime of our interest. In the following, we present the results.
@@ III Scaling Analysis at the AA Criticality
We start our study of the behavior of the scaling exponents right at the AA criticality, i.e., $\delta=0$. In a finite system, at this point, the wavefunction of the particle is neither localized nor delocalized, rather there happens to be a continuous phase transition. To quantify the localized or delocalized nature of the wavefunction, we use the observable localization length, $\zeta$, defined as,
@@
where $i$ denotes the lattice site, $p_{i}$ is the single-particle probability density at site $i$ and $i_{c}$ is the localization center having the expression $i_{c}=\sum_{i=1}^{L}ip_{i}$. Expanding the $n^{\mathrm{th}}$ normalised eigenstate of the system, $|\psi^{n}\rangle$, in terms of the single-particle computational basis, $|i\rangle$, such that $|\psi_{n}\rangle=\sum_{i}c^{(i)}_{n}|i\rangle$, $p_{i}$ is given by $p_{i}=|\langle i|\psi^{n}\rangle|^{2}=|c^{(i)}_{n}|^{2}$. If $g$ is the control parameter in a certain localization-delocalization transition and $g_{c}$ is the critical point then in the thermodynamic limit, near this criticality, the localization length scales as,
@@
Figure 2: Scaling analysis at $\delta=0$: (a1) presents localization length $\zeta$ of the ground state against Stark field strength $h$ for different system sizes, $L=144$ (blue circle), $L=233$ (violet square), $L=377$ (indigo diamond), $L=610$ (maroon triangle up) and $L=987$ (orange triangle left). (a2) shows a collapse plot for the localization length with the value of scaling exponent $\nu=0.29$. (b1) displays the IPR versus $h$ for various values of $L$ which is aforementioned in (a1). (b2) demonstrates the collapse plot of IPR using the value of scaling exponent $s=0.096$ and $\nu=0.29$. (c1) shows the energy gap $\Delta E$ between the ground state and the first excited state as a function of $h$ for different $L$ values as mentioned previously. (c2) demonstrates the collapse plot of $\Delta E$ using the scaling exponent $z=2.37$. we have taken $\phi$ average for 5000 random choices in the range $[0,1]$.
@@
where $f_{1}[\cdot]$ is an arbitrary function. This scaling ansatz is confirmed through the methodology of data collapse. The concept behind data collapse is as follows: We collect the data $\zeta/L$ versus $hL^{1/\nu}$ for various system sizes $L$, with a specific value of $\nu$, for which all the datasets collapse perfectly. Initially, by some trial and error, we manually find out a rough range of the scaling exponent for which the data collapse is quite well. Furthermore, for a better estimation of the scaling exponent we employ the cost function approach (see Appendix A for details).
Given that the details of $\zeta$ depend on a particular choice of $\phi$, a configuration averaging over 5000 random realizations of $\phi$ is performed to eliminate the $\phi$ dependent fluctuations. In the single-particle case, apart from $\zeta$, similar averaging is performed for all other quantities of our interest, to be introduced later. In Fig. 2(a1), we depicted $\zeta$ against $h$ for various values of $L$. This figure illustrates that for finite-sized systems, an initial flat region exists wherein the corresponding wave function exhibits an extended nature, indicating that the ground state of the Hamiltonian for $\delta=0$ is equally probable across this phase. Beyond a certain threshold of $h$, the localization length exhibits weak dependence on the system size, leading to the localization of the wave function.
Setting $|g-g_{c}|=h$ in Eq. (3), one finds the localization length exponent to follow the relation $\zeta\propto h^{-\nu}$, where the scaling exponent $\nu$ controls the rate of divergence of the localization length of the system in the thermodynamic limit near criticality. The scaling exponent, $\nu$, can be computed via two explicit approaches. A straightforward approach is to perform a simple fitting analysis on the finite-size data set. This is what is shown via the green dotted line in Fig. 2(a1). The best fit suggests $\nu=0.296(2)$, where (.)(.)( . ) is the fitting error on the last significant digit reported in this work. Hence, we obtain $\nu$ within $0.6\%$ uncertainty via best-fit. The data collapse method provides a second route for extracting $\nu$. Here the scaled localization length, $\zeta/L$, is plotted with the scaled Stark field, $hL^{1/\nu}$. Finite size data collapse can be obtained by tuning $\nu$ (see Eq. (5)). The quality of collapse is then examined via computation of the cost function, $C_{Q}$ (see Appendix A). The best collapse corresponds to the particular $\nu$, for which $C_{Q}$ is the minimum. The cost function approach further suggests the scaling exponent $\nu$ within a window $(\Delta\nu)$ between $0.286$ to $0.298$. $C_{Q}$ assumes minimal value within this window and remains flat throughout (see Appendix A). We report the optimal value of $\nu$ by taking an average over the flat window of $\Delta\nu$ and the uncertainty associated with the estimated value of the quantity is ±$\pm\Delta\nu/2$. Hence, the optimal $\nu$ corresponding to the best collapse turns out to be $0.292(6)$. We demonstrate a collapse plot in Fig. 2(a2) utilizing the function specified previously in Eq. (5) for $\nu=0.292$. Evidently, the reported values of $\nu$ from two different methods agree within the uncertainty range. We perform similar error analysis for all the cases to be discussed in the following.
@@
The variation in the value of IPR with the increase in $h$ for different system sizes $L$ is depicted in Fig. 2(b1). In this figure, up to a certain region of $h$, the IPR values appear flat, indicating the delocalized nature of the wavefunction. Beyond this region, the IPR becomes independent of the system size and hence suggests the system’s entrance into the localized phase. Considering that the IPR scales with $h$ as, $\text{IPR}\sim h^{s}$, green-colored dashed line has corresponds to the best fit. The obtained value of $s$ from the best fit is given by, $s=0.097(2)$. Moreover, we again perform the cost function analysis for IPR as well. It suggests the value of $s$ as, $s=0.0968(1)$. Fig. 2(b2) presents the best collapse plot of the IPR for $s=0.0968$. Noticeably, in this case, the ratio $s/\nu$ $(s/\nu\approx 0.33)$ coincides with the one from the pure AA model. Hence, the scaling of IPR with the system size is basically the same for both of these two models.
@@
Fig. 2(c1) shows, the energy gap $\Delta E$ against the Stark field strength $h$ for different system sizes. In this case, also there is an initial flat region showing a very narrow energy gap in the delocalized phase. Beyond a certain strength of Stark field $h$, $\Delta E$ becomes independent of the system size, and the energy gap gets increased with an increase in the field strength. From the fitting function $\Delta E\sim h^{\nu z}$, as depicted by the green-colored dashed line in Fig. 2(c1), We obtain, $\nu z=0.702(1)$. The value of $z$ can be derived by simply dividing $\nu z$ by $\nu$. We find $z=2.371(1)$. The cost function approach via data collapse also suggests that $z=2.371(1)$. Similar to the scaling exponent associated with IPR, the scaling exponent $z$ again remains the same as the pure AA model. The collapse plot corresponding to $\Delta E$ is presented in Fig. 2(c2).